I have a question regarding a previous question that was answered. It's under question #144072:

Q:Solve by factoring.
The radius of a circle is increased by 3 in., which increases the area by 100 in^2. Find the radius of the original circle. Round to the nearest hundredth.
Please answer.This question is from textbook Beginning Algebra with Applications

The answer was given as:
1. pi*r^2=A original
(this is clear)
2. pi*(r+3)^2=A+100 new
(this is clear)
3. pi(r^2+6r+9)=A+100
(this is clear)
4. pi(r^2+6r+9)-100=A
(this is clear)
5. pi(r^2+6r+9)-100=pi*r^2
(please explain this step? I don't understand how the transition took place from moving the "pi" from the front to placing it behind the 100 in the parentheses)? Please write out the step?
6. r^2+6r+9-(100/pi)=r^2

(this is clear)
7. 6r+9-31.83=0
(this is clear)
8. 6r=22.83
(this is clear)
9. r=3.81






Answer by Alan3354(69443)

5. pi(r^2+6r+9)-100=pi*r^2
(please explain this step? I don't understand how the transition took place from moving the "pi" from the front to placing it behind the 100 in the parentheses)? Please write out the step?
-----------------
The left side is the area after r is increased by 3 to (r+3). Subtracting the 100 makes it equal to the area when the radius was still r.
------------------
6. r^2+6r+9-(100/pi)=r^2
(this is clear)
7. 6r+9-31.83=0
(this is clear)
8. 6r=22.83
(this is clear)
9. r=3.81